OpenStudy (anonymous):

Which of the following is an equation of the line that passes through (3,-5) and has the slope of 2. y+5=2(x-3),y+5=2(x+3),y-5=3(x-2),y+5=3(x+2) ?

OpenStudy (kira_yamato):

Using the formula y - y1 = m (x - x1), y + 5 = 2 (x - 3)